I need to solve......144x^2 + 64 = 0....I do not even remember where to begin.
Answer by Alan3354(69443)

144x^2 + 64 = 0
9x^2 + 4 = 0
9x^2 = -4
x^2 = -4/9
x = ± 2i/3 (i = sqrt(-1) )